OpenStudy (anonymous):

Write the quadratic equation in factored form. Be sure to write the entire equation.

OpenStudy (solomonzelman):

more info maybe ?

OpenStudy (anonymous):

x 2 + x - 12 = 0

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm x^2+x-12=0 }\) \(\large\color{midnightblue}{ \rm (-3) \times (4) =-12 }\) \(\large\color{midnightblue}{ \rm (-3)+ (4) =1 }\)

OpenStudy (solomonzelman):

So how would you factor it ?